OpenStudy (anonymous):

f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

someone help me i need help big time plz

OpenStudy (anonymous):

okay so f(5)=f(1) + f(2) + f(5- 1)

OpenStudy (anonymous):

we need to know what f(4) is in there right

OpenStudy (anonymous):

f(5)=f(1)+f(2)+f(4) f(4)=f(1)+f(2)+f(3) f(3)=f(1)+f(2)+f(2)<-- now this we know

OpenStudy (anonymous):

so work from bottom to up... do you get it?